# Coupled Variational Reinforcement Learning for Language Model General Reasoning

### v1

@@ Abstract
While reinforcement learning have achieved impressive progress in language model reasoning, they are constrained by the requirement for verifiable rewards. Recent verifier-free RL methods address this limitation by utilizing the intrinsic probabilities of LLMs generating reference answers as reward signals. However, these approaches typically sample reasoning traces conditioned only on the question. This design decouples reasoning-trace sampling from answer information, leading to inefficient exploration and incoherence between traces and final answers. In this paper, we propose C̱oupled V̱ariational Ṟeinforcement Ḻearning (CoVRL), which bridges variational inference and reinforcement learning by coupling prior and posterior distributions through a hybrid sampling strategy. By constructing and optimizing a composite distribution that integrates these two distributions, CoVRL enables efficient exploration while preserving strong thought-answer coherence. Extensive experiments on mathematical and general reasoning benchmarks show that CoVRL improves performance by 12.4% over the base model and achieves an additional 2.3% improvement over strong state-of-the-art verifier-free RL baselines, providing a principled framework for enhancing the general reasoning capabilities of language models.
@@ 1 Introduction
Recent works (deepseekai2025deepseekr1incentivizingreasoningcapability; yue2025vapoefficientreliablereinforcement; yu2025dapoopensourcellmreinforcement) explore reinforcement learning with verifiable rewards (RLVR) to enhance LLMs’ reasoning capabilities, demonstrating impressive results in mathematical tasks. In this paradigm, the LLM generates a chain of thought (wei2023chainofthoughtpromptingelicitsreasoning) followed by an answer, and a rule-based program evaluates the final answer, assigning binary rewards based on correctness. The model is then optimized using policy gradient methods such as GRPO (shao2024deepseekmathpushinglimitsmathematical). However, these techniques require verifiable rewards, which restricts their applicability to other domains where formal verification of model answers is challenging or impossible.
Figure 1: Comparison between verifier-free RL with question-only training and CoVRL. Unlike prior methods that sample reasoning traces conditioned only on the question, CoVRL couples question-conditioned prior sampling with answer-conditioned posterior sampling via a hybrid variational framework, enabling efficient exploration while preserving strong trace–answer coherence.
A natural solution to this constraint is to introduce an LLM as a verifier (ma2025generalreasoneradvancingllmreasoning), which performs a role similar to that of the reward model in reinforcement learning from human feedback (ouyang2022traininglanguagemodelsfollow). However, this approach requires a strong verifier and faces the risk of reward hacking (skalse2025definingcharacterizingrewardhacking). As an alternative, verifier-free methods (yu2025rlprextrapolatingrlvrgeneral; tang2025verifiablerewardsscalingreinforcement; zhou2025reinforcinggeneralreasoningverifiers; chen2024languagemodelshiddenreasoners) have emerged that utilize the probabilities LLMs assign to correct answers as reward signals. These methods view answer generation as a process that relies on a set of implicit reasoning steps. In this formulation, each potential reasoning trace is treated as a latent variable contributing to the final answer. The probability of the correct answer is obtained by marginalizing over all plausible reasoning traces conditioned on the input question. In practice, the model is trained using reinforcement learning by sampling different intermediate thoughts for each question and using the probability of generating the correct answer from each thought as the reward.
@@
In this work, we propose Coupled Variational Reinforcement Learning (CoVRL), which frames reasoning training as a variational optimization problem that couples prior and posterior distributions. The core idea is to address the challenges of question-only training by establishing coupling between two complementary distributions for sampling reasoning traces during training. We leverage both prior and posterior distributions as variational components, corresponding to two complementary generation modes: (1) question-only generation, which samples traces from the prior distribution, reflecting real inference conditions; and (2) answer-guided generation, which samples traces from the posterior distribution to generate coherent reasoning that leads to the correct answer. This dual-mode strategy provides answer guidance during training while ensuring that the learned reasoning patterns transfer effectively to inference, thereby improving sample efficiency and mitigating incoherence between reasoning traces and answers.
@@ 2.1 Preliminary
Recent work (zhao2025geometricmeanpolicyoptimization; yu2025dapoopensourcellmreinforcement) has explored using reinforcement learning to optimize language models for reasoning tasks. Given a reward function $R(x,y)$ that evaluates the correctness of the generated reasoning and answer, the objective is:
@@
where $p_{\theta}$ denotes the model distribution parameterized by $\theta$, $x$ is the question, and $y$ represents the generated reasoning and answer. This objective can be optimized using policy gradient methods like GRPO (shao2024deepseekmathpushinglimitsmathematical). However, these approaches require verifiable reward signals, limiting their applicability to domains without accessible verifiers.
@@
Variational inference (li2019stablevariationalautoencodertext; kingma2022autoencodingvariationalbayes) provides a principled framework for learning models with latent variables. Consider a generative process where observed data $y$ depends on latent variables $z$:
@@
Here, $p_{\theta}(y|z)$ is the decoder distribution that reconstructs the observed data given the latent variables, while $q_{\phi}(z|y)$ is the encoder distribution that infers latent representations from the observed data. This lower bound is called the evidence lower bound (ELBO), which is widely used in variational autoencoders (kingma2022autoencodingvariationalbayes) and diffusion models (ho2020denoisingdiffusionprobabilisticmodels).
@@ 2.2 Coupled Variational Reinforcement Learning
Previous approaches can be viewed as setting $q(z)=p_{\phi}(z|x)$, which suffers from the aforementioned limitations. A natural alternative is to use the posterior distribution as $q(z)$, as in VAE (kingma2022autoencodingvariationalbayes), which samples reasoning traces conditioned on both the question and the target answer, i.e., $q(z)=q_{\psi}(z|x,y)$. This enables generating more relevant reasoning paths and improves sampling efficiency compared to prior-only sampling.
However, this introduces a fundamental training-inference mismatch. During training, we optimize the posterior distribution with access to target answers, but at inference time, we must rely on the prior distribution $p_{\phi}(z|x)$, which is conditioned only on the question. Even with KL divergence regularization $D_{\text{KL}}(q_{\psi}(z|x,y)\|p_{\phi}(z|x))$ that encourages the posterior to stay close to the prior, there is no guarantee that optimizing the posterior will improve the prior due to the asymmetric nature of reverse KL divergence (chan2022greedificationoperatorspolicyoptimization). Specifically, while $D_{\text{KL}}(q_{\psi}\|p_{\phi})$ constrains the posterior to avoid low-probability regions of the prior, it cannot ensure that the posterior covers all high-probability regions of the prior. This results in inadequate training for some regions of the prior, potentially leading to poor inference performance when sampling from these regions. In response to the difficulties of independently sampling from the prior $p_{\phi}(z|x)$ and posterior $q_{\psi}(z|x,y)$ distributions, we introduce CoVRL, which establishes coupling between these distributions through complementary mechanisms.
2.3 Composite Distribution and Coupled Optimization
We first construct a composite distribution that unifies both question-only and answer-guided generation modes:
$p^{\prime}(z|x,y)=\frac{1}{2}p_{\phi}(z|x)+\frac{1}{2}q_{\psi}(z|x,y)$ (6)
This composite distribution establishes coupling between the prior $p_{\phi}(z|x)$ and posterior $q_{\psi}(z|x,y)$ distributions. By replacing $q(z)$ with this composite distribution $p^{\prime}(z|x,y)$, we reformulate the ELBO from Equation 5 as:
@@ 2.3 Composite Distribution and Coupled Optimization
Reconstruction term
$\mathbb{E}_{p^{\prime}(z|x,y)}[\log p_{\theta}(y|z,x)]$ measures the model’s ability to predict the correct answer $y$ given the reasoning trace $z$ and question $x$, serving as the primary target for improving reasoning quality.
Regularization term
$D_{\text{KL}}(p^{\prime}(z|x,y)\|p_{\phi}(z|x))$ constrains the composite distribution to remain close to the prior distribution, ensuring that the learned reasoning patterns transfer effectively to inference.
This coupled optimization framework addresses the training-inference mismatch by simultaneously optimizing both the prior and posterior distributions toward the same objective through the composite distribution.
In practice, all three components can be implemented using a single LLM with different prompts: $p_{\phi}(z|x)$ conditions on the input question, $q_{\psi}(z|x,y)$ conditions on both question and answer, and $p_{\theta}(y|z,x)$ predicts the answer given the reasoning trace. However, directly sampling from the composite distribution $p^{\prime}(z|x,y)$ would require real-time mixing of prior and posterior distributions for generating each new token. This approach requires additional computation and complex modifications to prevalent LLM-RL frameworks (sheng2024hybridflow; hu2024openrlhf), which typically utilize efficient inference engines such as SGLang (sglang2024) and vLLM (kwon2023efficient) for rollout. Therefore, instead of sampling from the actual composite distribution, we adopt off-policy hybrid sampling. For each training sample, we randomly select from the prior with probability $\alpha$ or the posterior with probability $(1-\alpha)$:
This approach ensures that both distributions are sampled, enabling joint optimization of exploratory and answer-guided reasoning.
For the Composite distribution
$p^{\prime}(z|x,y)=\frac{1}{2}p_{\phi}(z|x)+\frac{1}{2}q_{\psi}(z|x,y)$, the gradient becomes:
Optimizing the composite distribution requires policy gradient methods due to the discrete sampling process. We employ Group Relative Policy Optimization (GRPO) (shao2024deepseekmathpushinglimitsmathematical) to optimize the reconstruction term with the following objective:
[…]
where $r(z)=\frac{p^{\prime}_{\text{new}}(z|x,y)}{p^{\prime}_{\text{old}}(z|x,y)}$ represents the probability ratio between the new and old policies, and $\epsilon$ is the clipping parameter. The advantage $\hat{A}=\log p_{\theta_{\text{old}}}(y|z,x)-\bar{R}$ is computed using group relative estimation, where $\bar{R}$ is the average reward across the batch. Importantly, the reward is evaluated based on the model’s own answer prediction probability, eliminating the need for external verifiers.
However, the hybrid sampling strategy introduces a fundamental difference between the hybrid sampling distribution and the target composite distribution. Nevertheless, we can still optimize the composite distribution through importance sampling. Specifically, we replace $p^{\prime}_{\text{old}}(z|x,y)$ in Equation 10 with the actual sampling distribution $p_{\text{hybrid}}(z|x,y)$. Thus, the importance ratio $r(z)$ becomes:
$r(z)=\frac{p^{\prime}(z|x,y)}{p_{\text{hybrid}}(z|x,y)}$ (11)
This formulation enables mathematically principled optimization of the target composite distribution.
Prior distribution
For $p_{\phi}(z|x)$, the gradient is:
$\nabla_{\phi}$ $\mathbb{E}_{p_{\phi}(z|x)}[\log p_{\theta}(y|z,x)]=$ (12)
An important observation is that optimizing the prior distribution alone corresponds to standard maximum likelihood estimation, which can be efficiently computed using negative log-likelihood loss. However, applying NLL loss to all reasoning traces is problematic, as some traces may be of low quality. Since our model also serves as a reward model that evaluates reasoning quality through $\log p_{\theta}(y|z,x)$, we selectively reinforce high-quality traces. Therefore, we only compute the loss on samples that complete successfully and have positive advantage $\hat{A}>0$.
@@
2.4 Optimizing KL Divergence
There are two approaches for optimizing the KL divergence term: (1) incorporating KL as part of the reward function as in PPO (schulman2017proximalpolicyoptimizationalgorithms), or (2) treating KL as a separate loss term as in GRPO (shao2024deepseekmathpushinglimitsmathematical). We adopt the latter approach, which we find is more stable.
Optimizing KL divergence as a loss function requires a non-negative and low-variance KL estimator. We extend existing KL estimators to handle our composite distribution setting, which couples prior and posterior through the KL regularization term. Specifically, we need to estimate $D_{\text{KL}}(p^{\prime}(z|x,y)\|p_{\phi}(z|x))$ where $p^{\prime}(z|x,y)=\frac{1}{2}p_{\phi}(z|x)+\frac{1}{2}q_{\psi}(z|x,y)$. We extend previous low-variance KL estimators (schulman2020kl) and derive different estimators depending on the sampling distribution using importance sampling with Bregman divergence-based control variates.
When sampling from the prior distribution:
When sampling from the posterior distribution:
@@ 2.5 Prompt Template and Tokenization
We follow the template format suggested in previous work (deepseekai2025deepseekr1incentivizingreasoningcapability), with simple extensions to suit our hybrid sampling strategy. The prompt templates for prior and posterior generation are shown in Figure 3. For tokenization, we separately encode special tokens including <think>, </think>, <answer> and </answer> to ensure consistent token ID sequences. However, during model generation, these special tokens may still be merged with adjacent tokens. Therefore, we use strings as stop signals for model generation, and after generation completes, we decode the outputs and re-encode them with separately encoded special tokens.
3 Experiment
3.1 Setup
Dataset
Following previous work, we use the dataset curated by (ma2025generalreasoneradvancingllmreasoning), which was sourced from WebInstruct (yue2024mammoth2scalinginstructionsweb). We retain only non-mathematical questions to assess training improvements on general reasoning abilities. Beyond this selection, we apply no additional filtering to evaluate the algorithm’s robustness across diverse question types, difficulty levels, and quality variations.
Implementation
We conduct experiments primarily using Qwen2.5-7B-Base (qwen2025qwen25technicalreport). We directly fine-tune the base model without an intermediate supervised fine-tuning stage. We implement our RL training pipeline using the verl framework (sheng2024hybridflow). During training, we use a batch size of 192 questions and generate 8 responses per question during rollout. For rollout sampling, we employ temperature=1.0, top_p=1.0, and max_tokens=2048 to ensure diverse response generation while maintaining reasonable response lengths. The clip threshold in the policy gradient loss is set to 0.3 given the off-policy nature of our algorithm. We use a cosine learning rate scheduler with 64 warmup steps and a peak learning rate of 1e-6.
Baselines
Our baseline selection focuses on other verifier-free methods that do not require external verification models during training: VeriFree (zhou2025reinforcinggeneralreasoningverifiers), RLPR (yu2025rlprextrapolatingrlvrgeneral), JLB (tang2025verifiablerewardsscalingreinforcement), RAVR (lin2025ravrreferenceanswerguidedvariationalreasoning), and LaTRO (chen2024languagemodelshiddenreasoners). We summarize the gradient estimators of these methods in Appendix B. To eliminate the influence of different RL techniques and focus on the gradient estimator, we train all baselines on the Qwen2.5-7B-base model using the GRPO (shao2024deepseekmathpushinglimitsmathematical) with identical hyperparameters on the filtered dataset.
Evaluation
We evaluate our method on a comprehensive suite of reasoning benchmarks. For general reasoning, we use MMLU-Pro (wang2024mmluprorobustchallengingmultitask), GPQA (rein2023gpqagraduatelevelgoogleproofqa), and TheoremQA (chen2023theoremqatheoremdrivenquestionanswering). For mathematical reasoning, we use AIME’24 (aime2024_i), AQuA (ling2017programinductionrationalegeneration), CARP-EN (zhang2023evaluatingimprovingtoolaugmentedcomputationintensive), MATH-500 (lightman2023lets), Minerva (lewkowycz2022solvingquantitativereasoningproblems), and SAT-Math (mcaleste_sat_dataset_2023). For evaluation, we use temperature=0.6 and max_tokens=4096. We utilize Math-Verify to check answer correctness; more details are provided in Appendix C. To reduce evaluation variance, we repeat the evaluation $N$ times for each dataset, where $N$ depends on the test set size. We report the average performance across these $N$ runs as Average@$N$ to account for sampling variance.
@@ 3.1 Setup
3.2 Main Results
Table 1 presents our main experimental results across all benchmarks. Our approach achieves consistent improvements over the base model and competitive performance against other verifier-free methods.
3.3 Training Dynamics
We present the evolution of key metrics in Figure 4 to better understand the algorithm. This reveals several observations:
The posterior distribution is effective in providing guidance.
Figure 4(a) shows steady reward score improvements throughout training, with the posterior distribution consistently outperforming the prior distribution. This persistent score gap validates our answer-guided sampling strategy and confirms that posterior sampling enables more efficient exploration of high-quality reasoning paths.
CoVRL enhances reasoning capabilities with prolonged chain-of-thought traces.
Figure 4(b) demonstrates a stable increase in response length, indicating that the model progressively generates more detailed reasoning processes. This trend suggests that CoVRL successfully encourages elaborate chain-of-thought reasoning with more thorough step-by-step explanations.
Regularization provides stable optimization dynamics.
Figures 4(c) and 4(d) show stable downward trends in NLL and KL losses. The decreasing losses indicate improved answer prediction and successful regularization, confirming that our variational objective effectively balances the reconstruction and regularization terms.
Figure 5: Impact of prior sampling probability $\alpha$ in hybrid sampling strategy.
3.4 Hybrid Sampling Strategy
We examine the impact of our hybrid sampling strategy by varying the mixing ratio $\alpha$ between prior $p_{\phi}(z|x)$ and posterior $q_{\psi}(z|x,y)$ distributions. Here $\alpha$ represents the probability of sampling from the prior distribution. By default, we set $\alpha=0.5$. Figure 5 demonstrates that low prior sampling probability ($\alpha=0.1$) outperforms high prior sampling probability ($\alpha=0.9$). This highlights the important role of the posterior distribution in our algorithm. When $\alpha=0.9$, the model primarily samples from the prior distribution, and we observe reduced reasoning chain length. We attribute this to difficulty in improving rewards, leading the model to prioritize minimizing KL loss and generating shorter sequences. When $\alpha=0.1$, reasoning chain length increases, and posterior-dominated sampling achieves better performance than prior-dominated sampling. However, due to training-inference mismatch, performance remains inferior to balanced sampling. Training dynamics under these settings are provided in Appendix D.

### v2

@@ Abstract
While reinforcement learning has achieved impressive progress in language model reasoning, it is constrained by the requirement for verifiable rewards. Recent verifier-free RL methods address this limitation by utilizing the probabilities that LLMs generate reference answers as reward signals. However, these approaches typically sample reasoning traces conditioned only on the question. This design decouples reasoning-trace sampling from answer information, leading to inefficient exploration and incoherence between traces and final answers. In this paper, we propose C̱oupled V̱ariational Ṟeinforcement Ḻearning (CoVRL), which bridges variational inference and reinforcement learning by coupling prior and posterior distributions through a hybrid sampling strategy. By constructing and optimizing a composite distribution that integrates these two distributions, CoVRL enables efficient exploration while preserving strong thought-answer coherence. Extensive experiments on mathematical and general reasoning benchmarks show that CoVRL improves performance by 12.4% over the base model and achieves an additional 2.3% improvement over state-of-the-art verifier-free RL baselines, providing a principled framework for enhancing the general reasoning capabilities of language models.
@@ 1 Introduction
Recent works (DeepSeek-AI et al., 2025; Yue et al., 2025; Yu et al., 2025a) explore reinforcement learning with verifiable rewards (RLVR) to enhance LLMs’ reasoning capabilities, demonstrating impressive results in mathematical tasks. In this paradigm, the LLM generates a chain of thought (Wei et al., 2023) followed by an answer, and a rule-based program evaluates the final answer, assigning binary rewards based on correctness. The model is then optimized using policy gradient methods such as GRPO (Shao et al., 2024). However, these techniques require verifiable rewards, which restricts their applicability to other domains where formal verification of model answers is challenging or impossible.
Figure 1: Comparison between verifier-free RL with question-only training and CoVRL. Unlike prior methods that sample reasoning traces conditioned only on the question, CoVRL couples question-conditioned prior sampling with answer-conditioned posterior sampling via a hybrid variational framework, enabling efficient exploration while preserving strong trace-answer coherence.
A natural solution to this constraint is to introduce an LLM as a verifier (Ma et al., 2025), which performs a role similar to that of the reward model in reinforcement learning from human feedback (Ouyang et al., 2022). However, this approach requires a strong verifier and faces the risk of reward hacking (Skalse et al., 2025). As an alternative, verifier-free methods (Yu et al., 2025b; Tang et al., 2025; Zhou et al., 2025a; Chen et al., 2024) have emerged that utilize the probabilities LLMs assign to correct answers as reward signals. These methods view answer generation as a process that relies on a set of implicit reasoning steps. In this formulation, each potential reasoning trace is treated as a latent variable contributing to the final answer. The probability of the correct answer is obtained by marginalizing over all plausible reasoning traces conditioned on the input question. In practice, the model is trained using reinforcement learning by sampling different intermediate thoughts for each question and using the probability of generating the correct answer from each thought as the reward.
@@
In this work, we propose Coupled Variational Reinforcement Learning (CoVRL), which frames reasoning training as a variational optimization problem that couples prior and posterior distributions. The core idea is to address the challenges of question-only training by establishing coupling between two complementary distributions for sampling reasoning traces during training. We leverage both prior and posterior distributions as variational components, corresponding to two complementary generation modes: (1) question-only generation, which samples traces from the prior distribution, reflecting real inference conditions; and (2) answer-guided generation, which samples traces from the posterior distribution to generate coherent reasoning that leads to the correct answer. This dual-mode strategy provides answer guidance during training while ensuring that the learned reasoning patterns transfer effectively to the inference phase, improving sample efficiency and mitigating incoherence between reasoning traces and answers.
@@ 2.1 Preliminaries
Recent work (Zhao et al., 2025; Yu et al., 2025a) has explored using reinforcement learning to optimize language models for reasoning tasks. Given a reward function $R(x,y)$ that evaluates the correctness of the generated reasoning and answer, the objective is:
@@
where $p_{\theta}$ denotes the model distribution parameterized by $\theta$, $x$ is the question, and $y$ represents the generated reasoning and answer. This objective can be optimized using policy gradient methods like GRPO (Shao et al., 2024). However, these approaches require verifiable reward signals, limiting their applicability to domains without accessible verifiers.
@@
Variational inference (Li et al., 2019; Kingma and Welling, 2022) provides a principled framework for learning models with latent variables. Consider a generative process where observed data $y$ depends on latent variables $z$:
@@
Here, $p_{\theta}(y|z)$ is the decoder distribution that reconstructs the observed data given the latent variables, while $q_{\phi}(z|y)$ is the encoder distribution that infers latent representations from the observed data. This lower bound is called the evidence lower bound (ELBO), which is widely used in variational autoencoders (Kingma and Welling, 2022) and diffusion models (Ho et al., 2020).
@@ 2.2 Coupled Variational Reinforcement Learning
Previous approaches can be viewed as setting $q(z)=p_{\phi}(z|x)$, which suffers from the aforementioned limitations. A natural alternative is to use the posterior distribution as $q(z)$, as in VAE (Kingma and Welling, 2022), which samples reasoning traces conditioned on both the question and the target answer, i.e., $q(z)=q_{\psi}(z|x,y)$. This enables generating more relevant reasoning paths and improves sampling efficiency compared to prior-only sampling.
However, this introduces a fundamental training-inference mismatch. During training, we optimize the posterior distribution with access to target answers, but at inference time, we must rely on the prior distribution $p_{\phi}(z|x)$, which is conditioned only on the question. Even with KL divergence regularization $D_{\text{KL}}(q_{\psi}(z|x,y)\|p_{\phi}(z|x))$ that encourages the posterior to stay close to the prior, there is no guarantee that optimizing the posterior will improve the prior due to the asymmetric nature of reverse KL divergence (Chan et al., 2022). Specifically, while $D_{\text{KL}}(q_{\psi}\|p_{\phi})$ constrains the posterior to avoid low-probability regions of the prior, it cannot ensure that the posterior covers all high-probability regions of the prior. This results in inadequate training for some regions of the prior, potentially leading to poor inference performance when sampling from these regions. In response to the difficulties of sampling independently from the prior $p_{\phi}(z|x)$ and posterior $q_{\psi}(z|x,y)$ distributions, we introduce CoVRL, which establishes coupling between these distributions through complementary mechanisms.
2.3 Composite Distribution
Rather than relying on either distribution independently, we propose to couple them through a unified framework. We first construct a composite distribution that mixes the prior and posterior at each token position. For a reasoning trace $z=(z_{1},\ldots,z_{T})$, the composite distribution is defined through token-level averaging:
$p^{\prime}(z_{t}|z_{<t},x,y)=\frac{1}{2}p_{\phi}(z_{t}|z_{<t},x)+\frac{1}{2}q_{\psi}(z_{t}|z_{<t},x,y)$ (6)
where the equal weighting ensures balanced gradient contributions from both the prior and posterior during optimization. The sequence-level probability is obtained by the standard autoregressive factorization:
$p^{\prime}(z|x,y)=\prod_{t=1}^{T}p^{\prime}(z_{t}|z_{<t},x,y)$ (7)
This composite distribution couples the prior $p_{\phi}(z|x)$ and posterior $q_{\psi}(z|x,y)$. By including both distributions in $p^{\prime}$, this construction enables direct optimization of the prior through answer-guided signals while ensuring the posterior aligns with inference conditions, avoiding the coverage issues of posterior-only training.
To derive the optimization objective for this composite distribution, we replace $q(z)$ with $p^{\prime}(z|x,y)$ in Equation 5:
@@ 2.3 Composite Distribution
Reconstruction term $\mathbb{E}_{p^{\prime}(z|x,y)}[\log p_{\theta}(y|z,x)]$ measures the model’s ability to predict the correct answer $y$ given the reasoning trace $z$ and question $x$, serving as the primary target for improving reasoning quality.
Regularization term $D_{\text{KL}}(p^{\prime}(z|x,y)\|p_{\phi}(z|x))$ constrains the composite distribution to remain close to the prior distribution, ensuring that the learned reasoning patterns transfer effectively to inference.
In practice, all three components $p_{\phi}(z|x)$, $q_{\psi}(z|x,y)$, and $p_{\theta}(y|z,x)$ can be implemented using a single LLM with different prompts: $p_{\phi}(z|x)$ conditions on the input question, $q_{\psi}(z|x,y)$ conditions on both question and answer, and $p_{\theta}(y|z,x)$ predicts the answer given the reasoning trace. Thus, the subscripts $\phi$, $\psi$, $\theta$ denote different conditioning contexts rather than separate model parameters.
@@
2.4 Hybrid Sampling and Coupled Optimization
To optimize the composite distribution, a straightforward approach would be to directly sample from $p^{\prime}(z|x,y)$ during training. However, this would require real-time mixing of prior and posterior distributions for generating each new token, demanding additional computation and complex modifications to prevalent LLM-RL frameworks (Sheng et al., 2024; Hu et al., 2024), which typically utilize efficient inference engines such as SGLang (Team, 2024) and vLLM (Kwon et al., 2023) for rollout.
Therefore, instead of sampling from the actual composite distribution, we adopt off-policy hybrid sampling. For each training sample, we randomly sample from the prior with probability $\alpha$ or the posterior with probability $(1-\alpha)$:
This approach ensures that both distributions are sampled, enabling joint optimization of exploratory and answer-guided reasoning. Although $p_{\text{hybrid}}$ differs from the target composite distribution $p^{\prime}$, we show below that we can still optimize $p^{\prime}$ by appropriately reweighting samples through importance sampling.
For the composite distribution $p^{\prime}(z|x,y)$ defined through token-level mixing, the gradient becomes:
Optimizing the composite distribution requires policy gradient methods due to the discrete sampling process. We employ Group Relative Policy Optimization (GRPO) (Shao et al., 2024) to optimize the reconstruction term with the following objective:
[…]
where $r_{t}=\frac{p^{\prime}_{\text{new}}(z_{t}|z_{<t},x,y)}{p^{\prime}_{\text{old}}(z_{t}|z_{<t},x,y)}$ represents the token-level probability ratio between the new and old policies, and $\epsilon$ is the clipping parameter. The advantage $\hat{A}=\log p_{\theta_{\text{old}}}(y|z,x)-\bar{R}$ is computed using group relative estimation, where $\bar{R}$ is the average reward across the batch. Importantly, the reward is evaluated based on the model’s own answer prediction probability, eliminating the need for external verifiers.
To account for the distribution mismatch between our hybrid sampling strategy and the target composite distribution, we optimize the composite distribution through importance sampling. Specifically, we replace $p^{\prime}_{\text{old}}(z_{t}|z_{<t},x,y)$ in Equation 11 with the actual sampling distribution $p_{\text{hybrid}}(z_{t}|z_{<t},x,y)$ to enable mathematically principled optimization of the target composite distribution. Thus, the token-level importance ratio $r_{t}$ becomes:
$r_{t}=\frac{p^{\prime}_{\text{new}}(z_{t}|z_{<t},x,y)}{p_{\text{hybrid}}(z_{t}|z_{<t},x,y)}$ (12)
where $p_{\text{hybrid}}(z_{t}|z_{<t},x,y)$ equals $p_{\phi}(z_{t}|z_{<t},x)$ for samples drawn from the prior distribution, and $q_{\psi}(z_{t}|z_{<t},x,y)$ for samples drawn from the posterior distribution. In practice, this importance ratio can be computed via two forward passes for each sampled reasoning trace $z$: one with the prior template to obtain $p_{\phi}(z|x)$ and one with the posterior template to obtain $q_{\psi}(z|x,y)$.
For the prior distribution $p_{\phi}(z|x)$, the gradient is:
$\nabla_{\phi}$ $\mathbb{E}_{p_{\phi}(z|x)}[\log p_{\theta}(y|z,x)]=$ (13)
An important observation is that optimizing the prior distribution alone corresponds to standard maximum likelihood estimation, which can be efficiently computed using negative log-likelihood loss. However, applying NLL loss to all reasoning traces is problematic, as some traces may be of low quality. Since our model also serves as a reward model that evaluates reasoning quality through $\log p_{\theta}(y|z,x)$, we selectively reinforce high-quality traces. Therefore, we only compute the loss on samples that produce valid formatted outputs and have positive advantage $\hat{A}>0$.
@@ 2.4 Hybrid Sampling and Coupled Optimization
2.5 Optimizing KL Divergence
There are two approaches for optimizing the KL divergence term: (1) incorporating KL as part of the reward function as in PPO (Schulman et al., 2017), or (2) treating KL as a separate loss term as in GRPO (Shao et al., 2024). We adopt the latter approach, which we find is more stable.
Optimizing KL divergence as a loss function requires a non-negative and low-variance KL estimator. To estimate $D_{\text{KL}}(p^{\prime}(z|x,y)\|p_{\phi}(z|x))$ in our composite distribution setting, we extend the low-variance estimators of Schulman (2020). Specifically, we derive different forms depending on the sampling distribution using importance sampling with Bregman divergence-based control variates.
When sampling from the prior distribution:
When sampling from the posterior distribution:
@@ 2.6 Prompt Template and Tokenization
We follow the template format suggested in previous work (DeepSeek-AI et al., 2025), with simple extensions to suit our hybrid sampling strategy. The prompt templates for prior and posterior generation are shown in Figure 3. For tokenization, we separately encode special tokens including <think>, </think>, <answer> and </answer> to ensure consistent token ID sequences. However, during model generation, these special tokens may still be merged with adjacent tokens. Therefore, we use strings as stop signals for model generation, and after generation completes, we decode the outputs and re-encode them with separately encoded special tokens.
@@
3 Experiment
3.1 Setup
Dataset. Following previous work, we use the dataset curated by (Ma et al., 2025), which was sourced from WebInstruct (Yue et al., 2024). We retain only the non-mathematical questions to assess training improvements on general reasoning abilities. Beyond this selection, we apply no additional filtering to evaluate the algorithm’s robustness across diverse question types, difficulty levels, and quality variations.
Implementation. We conduct experiments primarily using Qwen2.5-7B-Base (Qwen et al., 2025). We directly fine-tune the base model without an intermediate supervised fine-tuning stage. We implement our RL training pipeline using the verl framework (Sheng et al., 2024). During training, we use a batch size of 192 questions and generate 8 responses per question during rollout. For rollout sampling, we employ temperature=1.0, top_p=1.0, and max_tokens=2048 to ensure diverse response generation while maintaining reasonable response lengths. The clip threshold in the policy gradient loss is set to 0.3 given the off-policy nature of our algorithm. We use a cosine learning rate scheduler with 64 warmup steps and a peak learning rate of 1e-6.
Baselines. Our baseline selection focuses on other verifier-free methods that do not require external verification models during training: VeriFree (Zhou et al., 2025a), RLPR (Yu et al., 2025b), JLB (Tang et al., 2025), RAVR (Lin et al., 2025), and LaTRO (Chen et al., 2024). We summarize the gradient estimators of these methods in Appendix B. To eliminate the influence of different RL techniques and focus on the gradient estimator, we train all baselines on the Qwen2.5-7B-base model using GRPO (Shao et al., 2024) with identical hyperparameters on the filtered dataset.
Evaluation. We evaluate our method on a comprehensive suite of reasoning benchmarks. For general reasoning, we use MMLU-Pro (Wang et al., 2024), GPQA (Rein et al., 2023), and TheoremQA (Chen et al., 2023). For mathematical reasoning, we use AIME’24 (Mathematical Association of America, 2024), AQuA (Ling et al., 2017), CARP-EN (Zhang et al., 2023), MATH-500 (Lightman et al., 2023), Minerva (Lewkowycz et al., 2022), and SAT-Math (mcaleste, 2023). For evaluation, we use temperature=0.6 and max_tokens=4096. We utilize Math-Verify to check answer correctness; more details are provided in Appendix C. To reduce evaluation variance, we repeat the evaluation $N$ times for each dataset, where $N$ depends on the test set size. We report the average performance across these $N$ runs as Average@$N$ to account for sampling variance.
Figure 5: Impact of hybrid sampling probability $\alpha$.